Rudy Sandberg wants to invest in four-year bonds that are currently priced at $863.10. These bonds have a coupon rate of 6.37 percent and make semiannual coupon payments. What is the current market yield on this bond?
Answer #1

Formula : Current market semi-annual yield on the bond = Semi-annual coupon / Current market price

Semi-annual coupon = (1000 * 6.37%) / 2 = $63.70 / 2 = $31.85

Current market price of the bond = $863.10

Current market semi-annual yield on the bond = $31.85 / $863.10 = 0.0369 or 3.69%

Current market annual yield on the bond = 3.69% * 2  = 0.0738 or 7.38%